OpenStudy (anonymous):

what is the vertex of the graph of the function y=|-3x+2|-4 Options A-(-2/3,-4) B-(2/3,-4 C-(2/3,4) D-(-2/3,4) PLEASE SHOW WORK PLEASE HELP ME

jimthompson5910 (jim_thompson5910):

y = |-3x+2|-4 y = |-3(x-2/3)|-4 ... factor out -3, we want x to have a coefficient of 1 y = |-3|*|x-2/3|-4 y = 3|x-2/3|-4 The equation is now in the form y = a|x-h| + k The vertex of y = a|x-h| + k is (h,k) See what to do next?

OpenStudy (anonymous):

i dont kno because i never learned this that is y im asking for help

jimthompson5910 (jim_thompson5910):

compare y = a|x-h| + k with y = 3|x-2/3|-4 Notice how a = 3 What is the value of h? What is the value of k?

OpenStudy (anonymous):

k is -4 and h is -2/3

jimthompson5910 (jim_thompson5910):

close, k = -4 is correct

jimthompson5910 (jim_thompson5910):

but h is positive

jimthompson5910 (jim_thompson5910):

x-h x-2/3 so h = 2/3

jimthompson5910 (jim_thompson5910):

The vertex is then (h,k) So the vertex in this case is (2/3, -4)

OpenStudy (anonymous):

ok i get it now

jimthompson5910 (jim_thompson5910):

ok great